One-year Treasury securities yield 1 percent, 2-year Treasury securities yield 2.8 percent, and 3-year Treasury securities yield 5.5 percent. Assume that the expectations theory holds. What does the market expect will be the yield on 1-year Treasury securities two years from now?
Answer #1

Using Pure Expectation Theory,

(1 + r1,3)3 = (1 + r1,2)2(1 + r3,1)

(1 + 0.055)3 = (1 + 0.028)2(1 + r3,1)

r3,1 = 11.11%